# SAVGBench: Benchmarking Spatially Aligned Audio-Video Generation

### v1

@@ 5 Experimental Evaluation
In our experiments, we use the development set of the SVGSA24 dataset to train the Stereo MM-Diffusion model with its super-resolution model. The Stereo MM-Diffusion is trained from scratch while the super-resolution model is initialized with a pretrained model from the guided diffusion model [4], which is pretrained on ImageNet. To finetune the super-resolution model, we extract all videos as frames and train the model on each frame. We set a batch size to 4 with 8 NVIDIA A100 GPUs to train both models. In evaluation, we use the evaluation set of SVGSA24 as a reference to measure the quality of generated models.
We evaluate the quality and spatial alignment as shown in Table 1. The MM-Diffusion baseline achieves the Spatial AV-Align metric 0.48. This result demonstrates that there is room for improvement to increase the model’s ability.
Model FVD $\downarrow$ FAD $\downarrow$ Spatial AV-Align $\uparrow$
MM-Diffusion 1050.3 9.65 0.48
Ground Truth 572.05 3.70 0.92
Table 1: Evaluation results of the baseline and ground truth on key metrics. FVD and FAD evaluate the quality of video and audio, respectively, while our proposed Spatial AV Align metric indicates the alignment between both modalities.
@@ 6 Conclusion
This paper presents a new benchmark for Spatially Aligned Audio-Video Generation (SAVG). We propose three key components for the benchmark: dataset, baseline, and metrics. We introduce a spatially aligned audio-visual dataset named SVGSA24, which enables us to train and evaluate a model that generates videos with spatially aligned stereo audio. We also propose an audio-visual diffusion model focused on stereo audio-visual learning to accommodate spatial sound. We also introduce a new metric, Spatial AV-Align, which evaluates spatial alignment between audio and video using an object detector and a sound event localization and detection (SELD) model. Our experimental result shows that gaps exist between the baseline model and ground truth regarding video and audio quality, and spatial alignment between both modalities. This benchmark encourages future work in SAVG.
Appendix
Appendix A Implementation Details on Stereo SELD
We present the details of our stereo SELD model [1, 16]. Given stereo audio, two amplitude spectrograms and an inter-channel phase difference (IPD) spectrogram are extracted as input features for a neural network (NN)-based model. The model consists of convolution blocks, multi-head self-attention (MHSA) blocks, and fully connected layers. After receiving the features, the model estimates each class’s activity and horizontal position per frame. When a class’s activity exceeds a threshold, the SELD model outputs the class’s horizontal position. Since we consider onscreen sound events, the horizontal position ranges from 0 (left end of the corresponding image) to 1 (right end). The loss function for the activity is binary cross entropy. The loss function for the horizontal position is masked mean squared error (MSE), where we mask frames of no activity, i.e., we do not include them in loss computation. We train the model using stereo audio data extracted from the development set of the SVGSA24 dataset and evaluate it using the evaluation set. Note that the SVGSA24’s position labels are derived from the STARSS23 dataset. The fps of the SELD model is 10, the same as the fps of the position labels. In our experiments, the model achieved over 0.95 in each class’s F-score. We also found that the masked MSE model performed better than the MSE model.

### v2

@@ 5 Experimental Evaluation
5.1 Experimental Setup
In our experiments, we use the development set of the proposed dataset to train the baseline methods, including the super-resolution model. Note that the Stereo MM-Diffusion, video diffusion, and Stereo MMAudio models are trained from scratch. On the other hand, the super-resolution model is initialized with a pretrained model from the guided diffusion model [3], which is pretrained on ImageNet. To finetune the super-resolution model, we extract all videos as frames and train the model on each frame. Also, some components of Stereo MMAudio (i.e., VAE, Synchformer, and CLIP) are pretrained and kept frozen during training. We set a batch size of 32 to train all models. Further implementation details, such as the learning rate and number of iterations, are available in our GitHub repository. For evaluation, we use our proposed dataset’s evaluation set to assess model’s performance.
5.2 Experimental Result
We evaluate the quality and audio-visual alignment as shown in Table 1. We observe that the joint method achieves scores that are similar to or higher than those of the two-stage method. For distribution matching metrics (i.e., FVD, KVD, and FAD), the differences between the two approaches are minimal. Similarly, temporal AV-alignment scores are comparable. However, the joint method surpasses the two-stage method by a significant margin (approximately 0.15 points) in Spatial AV-Align scores. We hypothesize that this is because input conditions derived from generated videos are less plausible than those from ground truth videos and lack spatial information. In contrast, the joint method preserves spatial-temporal visual information during processing. Furthermore, compared to the Spatial AV-Align score on the evaluation set (i.e., ground truth), this result indicates potential for further improvement.
5.3 Subjective Test
Table 2: Subjective test results of the baseline results and ground truth data in terms of spatial audio-visual alignment.
Method MOS spatial audio-visual alignment $\uparrow$
Ground Truth 4.06
Joint Method 2.90
Two-stage Method 2.44
We conduct the mean opinion score (MOS) test on spatial audio-visual alignment with 7 participants, who are recruited in person and have expertise in audio-visual technologies. After passing a preliminary stereo listening test with a few samples using headphones, each participant then evaluates 45 randomly selected samples. Participants are asked to evaluate the spatial alignment of each audio-video sample on a 5-point Likert scale while listening through headphones, based on the question: “How would you rate the spatial alignment between the audio and video? (Focus on the direction of sounding sources and objects.)”
In this test, we report a trend similar to that seen when evaluating spatial alignment among ground truth videos, the joint method, and the two-stage method, as shown in Table 2. The joint method achieved a MOS of 2.90, whereas the two-stage method scored only 2.44 in this subjective listening evaluation. This result indicates that the joint method has better spatial alignment between visual and audio outputs compared to the two-stage method, following the trend in our Spatial AV-Align scores.
@@ 6 Conclusion
This paper presents a new benchmark for the Spatially Aligned Audio-Video Generation (SAVG) task. The SAVGBench comprises a well-curated audio-visual dataset and evaluation metrics on generated audio and video. The audio-visual dataset contains stereo audio and perspective video content that focuses on onscreen sound events, facilitating the training and evaluation of SAVG methods. The Spatial AV-Align metric assesses spatial alignment between audio and video by utilizing an object detector and a sound event localization and detection (SELD) model. It is applicable when both video and audio are generated, because ground truth audio is not required. Using the dataset and metrics, we benchmark two types of baseline methods: one is based on a joint audio-video generation model, and the other is a two-stage method that combines a video generation model with a video-to-audio generation model. Our experimental results indicate that gaps exist between the baseline methods and the ground truth in terms of video and audio quality, as well as spatial alignment between the two modalities. This benchmark encourages future work to fill the gaps, especially in spatial audio-visual alignment. Furthermore, other potential avenues include extending the benchmark to different indoor environments and more diverse sound events, and integrating text conditioning into the SAVG task.
7 Acknowledgments
We sincerely thank Ryosuke Sawata for his helpful feedback on this manuscript, Akio Hayakawa for his valuable advice on the MMAudio training and subjective test design, and Kengo Uchida for the constructive discussion about the evaluation metrics.